Explain this question please. I want to know how is my teacher solved it by table? I added my teacher table that he did to solve the problem.

H&T 180 $4 Heat capacity and specific heat Exercise 4.3 A piece of iron at 300°C with heat capacity 50 cal/C is placed in a tank of oil at 20°C with heat capacity 200 cal/C. Find the final imagining that heat is transferred from the iron to the oil until equilibrium is reached. You may imagine sending heat in any amounts you want; you do not have to use 1 calorie at a time. Keep track of your calculations in a table like the following: iToC) Temp, of iron (C) 300 Heat transferred Temp of o 0 cal 20 Check your results with a staff member.
media%2F133%2F13395825-d65b-4f1c-ab8e-69
Answer #1

Ans. The question wants us to keep track of the calculations in a table.

The specific heat of the iron = 50 cal/C. So when the iron loses 50 cal it's temperature decreases by 1 degree Celsius.

The specific heat of the oil = 200 cal.C, It means when the oil gets 200 cal , it's temperature rises by 1 degree Celsius.

The initial temperature of Iron = 300 C

Initial temperature of oil = 20 C

Now we decode the table provided.

1) Row 2: When 50 cal is transferred from iron to oil, the temperature of iron decreases by 1 C so it's at 299 C now, whereas the temperature of the oil increases by 50/200 = 0.25 C ( 200 cal is needed for 1 C rise), so the oil is at 20.25 C now.

2) Row 3: 5000 cal is transferred from the iron to the oil, so the iron's temperature decreases by 5000/50 = 100 C, so it's temperature is 300-100 = 200 C. The oil gains 5000 cal, so it's temperature increases by 5000/200 = 25C. Hence the temperature of the oil is 20 + 25 = 45 C

3) Row 4: 10000 cal is transferred from the iron to the oil, so the iron's temperature decreases by 10000/50 = 200 C, so it's temperature is 300-200 = 100 C. The oil gains 10000 cal, so it's temperature increases by 10000/200 = 50 C. Hence the temperature of the oil is 20. + 50 = 70 C

Similarly other two rows are obtained.